# Regularization implies balancedness in the Deep Linear Network

### v1

@@ 2.2. Differential geometry of $\mathcal{F}_{X}$
Let $F(\mathbf{W})=\|\mathbf{W}\|_{2}^{2}$. Then
$dF(\mathbf{W})\mathbf{w}_{\mathbf{a}}=\sum_{k=1}^{N-1}\mathrm{Tr}\left(G_{k}(a_{k}+a_{k}^{*})\right).$ (2.5)
@@
$F(\mathbf{W}(\tau))=\sum_{k=1}^{N}\mathrm{Tr}\left(W_{k}^{*}(\tau)W_{k}(\tau)\right)$
@@
$dF(\mathbf{W})\mathbf{w}_{\mathbf{a}}=-\mathrm{Tr}\left(a_{N-1}^{*}W_{N}^{*}W_{N}+W_{N}^{*}W_{N}a_{N-1}\right)$ (2.6)
@@ 2.3. Riemannian geometry of $\mathcal{F}_{X}$
$\langle\mathbf{w}_{\mathbf{b}},\mathbf{w}_{\mathbf{a}}\rangle=\mathrm{Tr}\left(b_{N-1}^{*}W_{N}^{*}W_{N}a_{N-1}\right)+$
$\sum_{k=2}^{N-1}\mathrm{Tr}\left((W_{k}^{*}b_{k}^{*}-b_{k-1}^{*}W_{k}^{*})(a_{k}W_{k}-W_{k}a_{k-1})\right)$
$+\mathrm{Tr}\left(W_{1}W_{1}^{*}b_{1}^{*}a_{1}\right).$
@@
$\langle\mathbf{w}_{\mathbf{b}},\mathbf{w}_{\mathbf{a}}\rangle=\sum_{k=1}^{N-1}\mathrm{Tr}(H_{k}(\mathbf{b})^{*}a_{k})=\sum_{k=1}^{N-1}\mathrm{Tr}(b_{k}^{*}H_{k}(\mathbf{a})).$ (2.9)
@@ 2.4. The regularizing flow
Let us now specialize to the case where $F(\mathbf{W})=\|\mathbf{W}\|_{2}^{2}$ is the $L^{2}$ regularizer. We now use Lemma 2 and Lemma 3 to obtain
@@
$\mathrm{grad}\|\mathbf{W}\|^{2}=\mathbf{w}_{\mathbf{b}}$ where $\mathbf{b}$ solves the block tridiagonal system
$H_{k}(\mathbf{b})+H_{k}^{*}(\mathbf{b})=4G_{k},\quad 1\leq k\leq N-1,$ (2.19)
along with the symmetry condition
@@
The block tridiagonal structure is more explicit when we use Definition 8 to see that equation (2.19) is equivalent to the system
@@
$-W_{k+1}^{*}(b_{k+1}+b_{k+1}^{*})W_{k+1}=4G_{k},\quad 1\leq k\leq N-1.$
Observe also that equation (2.19) further simplifies to $H_{k}=2G_{k}$ under the symmetry condition (2.20). We have written it as above to emphasize the manner in which one must solve for $\mathbf{b}$ given $\mathbf{G}$. Further, the definition of $H_{k}$ in equation (2.9) does not imply that $H_{k}=H_{k}^{*}$ in general. The symmetry condition is specific to the gradient flow of $\|\mathbf{W}\|^{2}$.
@@
$\mathrm{Tr}(H_{k}^{*}a_{k})=\mathrm{Tr}\left(G_{k}(a_{k}+a_{k}^{*})\right)=0,\quad 1\leq k\leq N-1.$ (2.22)
This identity holds for all $a_{k}\in\mathbb{M}_{d}$. It follows that $\mathrm{Tr}(H_{k}^{*}a_{k})=0$ when $a_{k}=-a_{k}^{*}$. Thus, $H_{k}=H_{k}^{*}$ since the space of Hermitian and anti-Hermitian matrices are orthogonal under the inner-product on $\mathbb{M}_{d}$ given by $\mathrm{Tr}$. But then we also have the identity $\mathrm{Tr}\left((H_{k}-2G_{k})a_{k}\right)=0$ for all Hermitian $a_{k}$. Since $H_{k}-2G_{k}$ is Hermitian, it follows that $H_{k}=2G_{k}$. ∎
@@
By Lemma 5, $\mathrm{grad}\|\mathbf{W}\|^{2}=\mathbf{w}_{\mathbf{b}}$ where $\mathbf{b}$ satisfies equation (2.19). Therefore, by Lemma 4 and Lemma 5
$\dot{\mathbf{G}}=d\mathbf{G}\mathbf{w}_{\mathbf{b}}=-4\mathbf{G}.$ (2.23)
@@
Our proof of Theorem 2 relies on explicit computations with the Riemannian manifold $(\mathcal{F}_{X},\iota)$. We present these calculations since they allow us to consider other gradient flows on $\mathcal{F}_{X}$, such as the Ness flow introduced below. However, the reader should note that the cancellations that lead to the closed form for $\dot{\mathbf{G}}$ have a simple geometric origin.
We first observe that the gradient of $\|\mathbf{W}\|_{2}^{2}$ in $\mathbb{M}_{d}^{N}$ is simply $2\mathbf{W}$. The gradient may then be decomposed into two components $\mathbf{w}_{\mathbf{b}}=\mathrm{grad}\|\mathbf{W}\|^{2}\in T_{\mathbf{W}}\mathcal{F}_{X}$ and $\mathbf{W}^{\perp}:=2\mathbf{W}-\mathbf{w}_{\mathbf{b}}\in T_{\mathbf{W}}\mathcal{F}_{X}^{\perp}$. The conservation laws for $\mathbf{G}$ are due to the fact that $T_{\mathbf{W}}\mathcal{F}_{X}^{\perp}$ lies in the nullspace of $d\mathbf{G}$. Thus,
$d\mathbf{G}\,\mathbf{w}_{\mathbf{b}}=d\mathbf{G}\,\mathbf{W}=-4\mathbf{G},$
@@
2.5. The Ness flow
The gradient flow of the squared moment map is an important tool in Ness’ work [31]. We derive its explicit form for the DLN by applying the calculations of Section 2.3 to the function
$\|\mathbf{G}\|_{2}^{2}=\sum_{k=1}^{N-1}\mathrm{Tr}(G_{k}^{*}G_{k})=\sum_{k=1}^{N-1}\mathrm{Tr}(G_{k}^{2}).$ (2.24)
@@ 2.5. The Ness flow
The gradient flow of $\|\mathbf{G}\|^{2}_{2}$ is expressed in coordinates as
$\dot{\mathbf{W}}=-2\mathbf{w}_{\mathbf{G}}.$ (2.26)
The corresponding evolution of the moments is given by
@@
For convenience of notation, let $F(\mathbf{W})=\|\mathbf{G}\|_{2}^{2}$. Then
$dF(\mathbf{W})\mathbf{w}_{\mathbf{a}}=2\sum_{k=1}^{N-1}\mathrm{Tr}\left(G_{k}dG_{k}\mathbf{w}_{\mathbf{a}}\right)=2\sum_{k=1}^{N-1}\mathrm{Tr}\left(G_{k}(H_{k}(\mathbf{a})+H_{k}(\mathbf{a})^{*}\right).$ (2.28)
@@
$\langle\mathrm{grad}F,\mathbf{w}_{\mathbf{a}}\rangle=\langle\mathbf{w}_{\mathbf{b}},\mathbf{w}_{\mathbf{a}}\rangle=\sum_{k=1}^{N-1}\mathrm{Tr}\left(b_{k}^{*}H_{k}(\mathbf{a})\right).$ (2.29)
@@
$\sum_{k=1}^{N-1}\mathrm{Tr}\left(b_{k}^{*}H_{k}(\mathbf{a})\right)=2\sum_{k=1}^{N-1}\mathrm{Tr}\left(G_{k}(H_{k}(\mathbf{a})+H_{k}(\mathbf{a})^{*}\right).$ (2.30)
When $X$ has full rank, $\mathbf{H}$ is an isomorphism. We may thus choose $\mathbf{a}$ such that $H_{k}=-H_{k}^{*}$, ensuring that $\mathrm{Tr}\left(b_{k}^{*}H_{k}(\mathbf{a})\right)=0$ whenever $H_{k}=-H_{k}^{*}$. Thus, $b_{k}=b_{k}^{*}$. It then also follows from equation (2.30) that $b_{k}=2G_{k}$. Equation (2.27) follows from Lemma 4 with $\mathbf{c}=2\mathbf{G}$. ∎
The Ness flow presents an interesting contrast with the regularizing flow. Lemma 5 shows that when we consider the functional $\|\mathbf{W}\|_{2}^{2}$, the gradient $\mathrm{grad}\|\mathbf{W}\|_{2}^{2}=\mathbf{w}_{\mathbf{b}}$ where $\mathbf{b}$ is given implicitly through the solution of the linear system  (2.19). This makes numerical implementations of the regularizing flow subtle, since one must solve for $\mathbf{b}$ at each step. However, despite the implicit nature of the regularizing flow, Theorem 2 tells that $\mathbf{G}$ evolves by pure scaling.
In contrast, $\mathrm{grad}\|\mathbf{G}\|_{2}^{2}=\mathbf{w}_{\mathbf{b}}$ where $\mathbf{b}=2\mathbf{G}$. Thus, the Ness flow (2.26) is explicit in $\mathbf{W}$ and does not require the solution of a linear system. On the other hand, while it is immediate from the definition of the gradient flow (2.26) that
$\frac{d}{dt}{\|\mathbf{G}\|_{2}^{2}}=-4\|\mathbf{w}_{\mathbf{G}}\|_{2}^{2},$ (2.31)
@@ 3.2. The Kempf-Ness theorem: abstract structure
We summarize the abstract setup following Helmke [14, §2]. The reader is also referred to [31] for finer results based on the gradient flow of the squared moment map.
@@
Ness uses the gradient flow of $\|\mu\|^{2}$ to classify the non-closed orbits, further stratifying them according to the minimal and non-minimal critical points of $\|\mu\|^{2}$ [31, Thm 6.2]). This analysis motivated our introduction of the Ness flow in Section 2.5.
@@ 3.3. Application of the Kempf-Ness theorem
$\mu(\mathbf{W})=2\mathbf{G}(\mathbf{W}).$ (3.7)
(This explains the factor of $2$ in several calculations, such as the proof of Lemma 5).
@@ 3.4. Hidden convexity
The Kempf-Ness theorem may be seen as an assertion that the squared norm function $\psi_{x}:\mathcal{G}\to\mathbb{R}$ has properties analogous to a convex function. In fact, the proof of the theorem begins with a consideration of ‘special functions’ on the line of the form $\sum_{i=1}a_{e}^{l_{i}x}$ where $a_{i}$ are positive numbers and the $l_{i}$ are arbitrary real numbers [20, §1]. Azad and Loeb noticed that the key feature of the squared norm function that is relevant to the Kempf-Ness theorem is its plurisubharmonicity, yielding the following
@@ 4. Acknowledgements
The authors express their gratitude to Yotam Alexander, Sanjeev Arora, Nadav Cohen, Colin McSwiggen, Linda Ness, Noam Razin, Karen Uhlenbeck and Tianmin Yu for stimulating conversations regarding this work.

### v2

@@ 2.2. Differential geometry of $\mathcal{F}_{X}$
Let $F(\mathbf{W})=\tfrac{1}{2}\|\mathbf{W}\|_{2}^{2}$. Then
$dF(\mathbf{W})\mathbf{w}_{\mathbf{a}}=\frac{1}{2}\sum_{k=1}^{N-1}\mathrm{Tr}\left(G_{k}(a_{k}+a_{k}^{*})\right)=\sum_{k=1}^{N-1}\mathrm{Re}\,\mathrm{Tr}\left(G_{k}^{*}a_{k}\right).$ (2.5)
@@
$2F(\mathbf{W}(\tau))=\sum_{k=1}^{N}\mathrm{Tr}\left(W_{k}^{*}(\tau)W_{k}(\tau)\right)$
@@
$2dF(\mathbf{W})\mathbf{w}_{\mathbf{a}}=-\mathrm{Tr}\left(a_{N-1}^{*}W_{N}^{*}W_{N}+W_{N}^{*}W_{N}a_{N-1}\right)$ (2.6)
@@ 2.3. Riemannian geometry of $\mathcal{F}_{X}$
$\langle\mathbf{w}_{\mathbf{b}},\mathbf{w}_{\mathbf{a}}\rangle=\mathrm{Re}\,\mathrm{Tr}\left(b_{N-1}^{*}W_{N}^{*}W_{N}a_{N-1}\right)+$
$\sum_{k=2}^{N-1}\mathrm{Re}\,\mathrm{Tr}\left((W_{k}^{*}b_{k}^{*}-b_{k-1}^{*}W_{k}^{*})(a_{k}W_{k}-W_{k}a_{k-1})\right)$
$+\mathrm{Re}\,\mathrm{Tr}\left(W_{1}W_{1}^{*}b_{1}^{*}a_{1}\right).$
@@
$\langle\mathbf{w}_{\mathbf{b}},\mathbf{w}_{\mathbf{a}}\rangle=\sum_{k=1}^{N-1}\mathrm{Re}\,\mathrm{Tr}(H_{k}(\mathbf{b})^{*}a_{k})=\sum_{k=1}^{N-1}\mathrm{Re}\,\mathrm{Tr}(b_{k}^{*}H_{k}(\mathbf{a})).$ (2.9)
@@ 2.4. The regularizing flow
Let us now specialize to the case where $F(\mathbf{W})=\tfrac{1}{2}\|\mathbf{W}\|_{2}^{2}$ is the $L^{2}$ regularizer. We now use Lemma 2 and Lemma 3 to obtain
@@
$\mathrm{grad}\tfrac{1}{2}\|\mathbf{W}\|^{2}=\mathbf{w}_{\mathbf{b}}$ where $\mathbf{b}$ solves the block tridiagonal system
$H_{k}(\mathbf{b})=G_{k},\quad 1\leq k\leq N-1.$ (2.19)
The block tridiagonal structure is as follows. Since $G_{k}=G_{k}^{*}$ we have
@@
We now use Definition 8 to see that equation (2.19) is equivalent to
@@
$-W_{k+1}^{*}(b_{k+1}+b_{k+1}^{*})W_{k+1}=2G_{k},\quad 1\leq k\leq N-1.$
@@
$\mathrm{Re}\mathrm{Tr}(H_{k}^{*}(\mathbf{b})a_{k})=\mathrm{Re}\mathrm{Tr}\left(G_{k}^{*}a_{k})\right)=0,\quad 1\leq k\leq N-1.$ (2.22)
This identity holds for all $a_{k}\in\mathbb{M}_{d}$. The theorem follows immediately because of the degeneracy of the inner-product defined in (1.14). ∎
@@
By Lemma 5, $\mathrm{grad}\tfrac{1}{2}\|\mathbf{W}\|^{2}=\mathbf{w}_{\mathbf{b}}$ where $\mathbf{b}$ satisfies equation (2.19). Therefore, by Lemma 4 and Lemma 5
$\dot{\mathbf{G}}=d\mathbf{G}\mathbf{w}_{\mathbf{b}}=-2\mathbf{G}.$ (2.23)
@@
Our proof of Theorem 2 relies on explicit computations with the Riemannian manifold $(\mathcal{F}_{X},\iota)$. We present these calculations since they allow us to consider other gradient flows on $\mathcal{F}_{X}$, such as the Kirwan-Ness flow introduced below. However, the reader should note that the cancellations that lead to the closed form for $\dot{\mathbf{G}}$ have a simple geometric origin.
We first observe that the gradient of $\tfrac{1}{2}\|\mathbf{W}\|_{2}^{2}$ in $\mathbb{M}_{d}^{N}$ is simply $\mathbf{W}$. The gradient may then be decomposed into two components $\mathbf{w}_{\mathbf{b}}=\mathrm{grad}\tfrac{1}{2}\|\mathbf{W}\|^{2}\in T_{\mathbf{W}}\mathcal{F}_{X}$ and $\mathbf{W}^{\perp}:=\mathbf{W}-\mathbf{w}_{\mathbf{b}}\in T_{\mathbf{W}}\mathcal{F}_{X}^{\perp}$. The conservation laws for $\mathbf{G}$ are due to the fact that $T_{\mathbf{W}}\mathcal{F}_{X}^{\perp}$ lies in the nullspace of $d\mathbf{G}$. Thus,
$d\mathbf{G}\,\mathbf{w}_{\mathbf{b}}=d\mathbf{G}\,\mathbf{W}=-2\mathbf{G},$
@@
2.5. The Kirwan-Ness flow
The Morse theory of the squared moment map is a fundamental concept in Kirwan and Ness’ work [25, 38]. We derive its gradient flow for the DLN by applying the calculations of Section 2.3 to the function
$\frac{1}{2}\|\mathbf{G}\|_{2}^{2}=\frac{1}{2}\sum_{k=1}^{N-1}\mathrm{Tr}(G_{k}^{*}G_{k})=\frac{1}{2}\sum_{k=1}^{N-1}\mathrm{Tr}(G_{k}^{2}).$ (2.24)
@@ 2.5. The Kirwan-Ness flow
The gradient flow of $\tfrac{1}{2}\|\mathbf{G}\|^{2}_{2}$ is
$\dot{\mathbf{W}}=-\mathbf{w}_{2\mathbf{G}},$ (2.26)
or equivalently, the cubic system (1.22). The corresponding evolution of the moments is given by
@@
For convenience of notation, let $F(\mathbf{W})=\tfrac{1}{2}\|\mathbf{G}\|_{2}^{2}$. Then
$dF(\mathbf{W})\mathbf{w}_{\mathbf{a}}=\sum_{k=1}^{N-1}\mathrm{Tr}\left(G_{k}dG_{k}\mathbf{w}_{\mathbf{a}}\right)=\sum_{k=1}^{N-1}\mathrm{Tr}\left(G_{k}(H_{k}(\mathbf{a})+H_{k}(\mathbf{a})^{*}\right).$ (2.28)
@@
$\langle\mathrm{grad}F,\mathbf{w}_{\mathbf{a}}\rangle=\langle\mathbf{w}_{\mathbf{b}},\mathbf{w}_{\mathbf{a}}\rangle=\sum_{k=1}^{N-1}\mathrm{Re}\mathrm{Tr}\left(b_{k}^{*}H_{k}(\mathbf{a})\right).$ (2.29)
@@
$\sum_{k=1}^{N-1}\mathrm{Re}\mathrm{Tr}\left(b_{k}^{*}H_{k}(\mathbf{a})\right)=\sum_{k=1}^{N-1}\mathrm{Tr}\left(G_{k}(H_{k}(\mathbf{a})+H_{k}(\mathbf{a})^{*}\right)=2\sum_{k=1}^{N-1}\mathrm{Re}\mathrm{Tr}\left(G_{k}(H_{k}(\mathbf{a})\right).$ (2.30)
When $X$ has full rank, $\mathbf{H}$ is an isomorphism. Thus, $b_{k}=2G_{k}^{*}=2G_{k}$. Equation (2.27) follows from Lemma 4 with $\mathbf{c}=2\mathbf{G}$. ∎
Remark 11.
While we assumed that $\mathbf{W}\in\mathcal{F}_{X}$ in order to define the gradient flow (2.26) we see that the flow is given by the cubic system (1.22) which is defined on all of $\mathbb{M}_{d}^{N}$. Thus, the flow may be easily implemented numerically. Its convergence as $t\to\infty$ follows from the general theory for the gradient flow of the squared moment map. An excellent exposition is provided by Lerman and adapting [29, Thm. 1] we have
Corollary 1.
The $\omega$-limit set $\omega(\mathbf{W})$ under the Kirwan-Ness flow for any $\mathbf{W}\in\mathbb{M}_{d}^{N}$ is a single point on the balanced variety $\mathcal{M}_{0}$.
The Kirwan-Ness flow presents an interesting contrast with the regularizing flow for $\tfrac{1}{2}\|\mathbf{W}\|^{2}$. Lemma 5 shows that when we consider the functional $\tfrac{1}{2}\|\mathbf{W}\|_{2}^{2}$, the gradient $\mathrm{grad}\tfrac{1}{2}\|\mathbf{W}\|_{2}^{2}=\mathbf{w}_{\mathbf{b}}$ where $\mathbf{b}$ is given implicitly through the solution of the linear system  (2.19). This makes numerical implementations of this regularizing flow subtle, since one must solve for $\mathbf{b}$ at each step. However, despite the implicit nature of the regularizing flow, Theorem 2 tells that $\mathbf{G}$ evolves by pure scaling.
In contrast, the Kirwan-Ness flow (2.26) is given explicitly by (1.22) and is easy to implement numerically. On the other hand, while it is immediate from the definition of the gradient flow (2.26) that
$\frac{d}{dt}\tfrac{1}{2}{\|\mathbf{G}\|_{2}^{2}}=-2\|\mathbf{w}_{\mathbf{G}}\|_{2}^{2},$ (2.31)
@@ 3.2. The Kempf-Ness theorem: abstract structure
We summarize the abstract setup following Helmke [18, §2]. The reader is also referred to [38] for finer results based on the gradient flow of the squared moment map.
@@
Ness used the gradient flow of $\|\mu\|^{2}$ to classify the non-closed orbits, further stratifying them according to the minimal and non-minimal critical points of $\|\mu\|^{2}$ [38, Thm 6.2]). This analysis motivated our introduction of the Kirwan-Ness flow in Section 2.5.
@@ 3.3. Application of the Kempf-Ness theorem
$\mu(\mathbf{W})=\mathbf{G}(\mathbf{W}).$ (3.7)
@@ 3.4. Hidden convexity
The Kempf-Ness theorem may be seen as an assertion that the squared norm function $\psi_{x}:\mathcal{G}\to\mathbb{R}$ has properties analogous to a convex function. In fact, the proof of the theorem begins with a consideration of ‘special functions’ on the line of the form $\sum_{i=1}a_{e}^{l_{i}x}$ where $a_{i}$ are positive numbers and the $l_{i}$ are arbitrary real numbers [24, §1]. Azad and Loeb noticed that the key feature of the squared norm function that is relevant to the Kempf-Ness theorem is its plurisubharmonicity, yielding the following
@@
4. Discussion: a new dynamic paradigm
4.1. Gradient flow of a regularized cost function
This is the gradient flow on $\mathbb{M}_{d}^{N}$
where the parameter $\kappa>0$ controls the strength of the regularization. The main observation then is that this dynamical system may be naturally decomposed at each point $\mathbf{W}$ into two orthogonal flows, one normal to $\mathcal{F}_{X}$ (learning) and the other parallel to $\mathcal{F}_{X}$ (regularizing). Our heuristic idea is that regularization is ‘fast’ because of the exponential rate of convergence provided by Theorem 2 (note that the rate is now $2\kappa$, not $2$), so that the dynamics of equation (4.1) may be rigorously approximated by the learning and regularizing flows.
4.2. Regularization by background noise
4.3. Is deep learning ‘secretly Bayesian’?
4.4. Conclusion
@@ 5. Acknowledgements
The authors express their gratitude to Yotam Alexander, Sanjeev Arora, Nadav Cohen, Boris Hanin, Colin McSwiggen, Linda Ness, Noam Razin, Karen Uhlenbeck and Tianmin Yu for stimulating conversations regarding this work.